A disoriented physics professor drives 3.25 km north, then 2.20 km west, and then 1.50 km south. Find the magnitude and direction of the resultant displacement using the component method.


1
Expert's answer
2021-10-11T08:55:09-0400

The displacement

"{\bf d}={\bf d_1}+{\bf d_2}+{\bf d_3}""d_x=d_{\rm East}=0+(-2.20)+0=-2.20\:\rm km""d_y=d_{\rm North}=3.25+0+(-1.50)=1.75\:\rm km"

Magnitude:

"d=\sqrt{d_x^2+d_y^2}=\sqrt{(-2.20)^2+1.75^2}=2.81\:\rm km"

Direction:

"\theta=\tan^{-1}\frac{d_y}{d_x}=\tan^{-1}\frac{1.75}{-2.20}=38.5^{\circ}\; \rm N\;of\;W"